OpenStudy (anonymous):

Doug has a weekend job. He earns $9.50 an hour and works 6 hours on Saturday. He works some more hours on Sunday, and his paycheck totals $133.00. How many hours did Doug work Sunday? A. 14 B. 10 C. 8 D. 6

OpenStudy (anonymous):

6*9.50=y where y is how much you made on saturday. Can you tell me how much that is? $57 the amount you make on sunday will be the total minus what you made on saturday $133.00-$57.00=$76.00 x*9.50=$76.00 divide both sides by 9.50 and you get your answer for x x is the number of hours. x=\frac{76.00}{9.50} 8

OpenStudy (anonymous):

\[x=\frac{ 76 }{ 9.5 }\]